And this one I'm really confused about. I gor the answer Fred is 14yo but it tells me the answer is wrong.
Fred is 3 years older than his sister Mary. In 7 years she will be six-sevenths of his age. How old is Fred?

Answer by stanbon(75887)

Fred is 3 years older than his sister Mary. In 7 years she will be six-sevenths of his age. How old is Fred?
---
Equations:
f = m + 3
(m+7) = (6/7)(f+7)
------------------------
Substitute for "m" and solve for "f":
f-3+7 = (6/7)(f+7)
-----
7(f+4) = 6(f+7)
7f + 28 = 6f + 42
f = 14 yrs (Fred's age now)
===============================
Note: You are correct.
===============================
Cheers,
Stan H.